VEGAS Team Meeting

February 04, 2013

Attendees

CALTECH, UC-B, & Columbia

Dan, Jayanth, Mark, Andrew, Glenn

NRAO:

Jason, Ray, Marty, Anish, Randy, Paul

Meeting Notes:

Next Teleconference: February 11, 2012 .

Postpone observations until the receiver is replaced.

Latest test results

  • additional FITS file fields should be added to capture counts and other useful troubleshooting items (D)

UTCdelta issue

UTCdelat issue is different now. In almost all files the first integration is large by about 5 micro sec. There are glitches in UTCdelta at about 5 nano sec (about one FPGA clock) (see utcglitch_jan27.png). In one of the files (/lustre/gbtdata/TGBT12B_504_06/VEGAS/2013_01_27_23:39:58) there is jump in utcdelta and the magnitude of the jump is different for different Banks (see utcd1_jan27.png).

Fixed.

Varying delays in bands
We see jumping of different times in varying bands of UTCdelta after an observation has run for a period (R)

Anish to re-run tests to attempt to replicate errors

Start time Alignment

Start time of data are aligned in almost all files. In two files, the start time of bank A is off by 28 micro sec.(D)

Re-test by doing several short iterations with existing function generator short gaps and see if they can be stitched together.

Step in time series

Noticed a step (~few percent) in the time series obtained from collapsing channels 100 to 900 (see timeser_jan27.png). May not be an issue for pulsar obs.

Re-test

Action Items from 28-Feb
  1. A BOF for mode 13, or some other mode that uses GPUs.
    • Not available yet
  2. A BOF with working switching signal logic.
    • designed but still troubleshooting switching table loading issues
  3. A test and decision on what computers to buy.
    • Not yet determined - Jayanth should go ahead with his tests to complete selection
  4. All Banks not starting at the same time.
    • (above)
  5. It will be good if we get the full spectral dynamic range corresponding to the 8 bit ADC in the next version of mode 1.
    • Conduct initial tests at Berkeley, to see if there are any limiting issues for dynamic range and optimize
  6. KFPA observation with mode 13.
    • awaits mode 13
  7. Manager/VEGAS HPC code that will let two instances of the program be run simultaneously.
    • not available, but not required for GC obs.
  8. Remember to do the RFI test of the rack.
    • Not yet scheduled
  9. Roach to Valon serial port connection.
    • Not yet implemented
  10. The I2C bus software tested with the manager and Galen's modules.
    • Post GC Obs
  11. The ROACH uboots reflased (1-6).
    • May be accomplished any time with access to system (Feb 5)
  12. The tcpborphserver init script needs configured.

VEGAS observation preparation Status:

  1. Sort out the timing issue in utcdelta
    • postpone obs until this is fixed
  2. Fix board crashes (maybe not a barrier to obs.)
  3. what tests should we do before we have enough confidence to start observing?
    • Test Pulsar
      • more tests after new fits fields are added

Action Items:

Other items:

-- MartyBloss - 2013-02-04
Topic revision: r2 - 2013-02-04, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback